OpenStudy (wcrmelissa2001):

Help on linear inequalities please! If x,y,z≥0, 3y+2z=3+x, 3y+z=4-3x, find the largest and smallest value of 3x-3y+4z. Every time I've done this I get a different answer. Thank yoU!

OpenStudy (phi):

do you know linear algebra? you can write the augmented matrix [ -1 3 2 3; 3 3 1 4] and find its reduced row echelon form [1 0 -¼ ¼; 0 1 7/12 13/12] which means x= ¼ z + ¼ y= -7/12 z +13/12 and 3x-3y+4z can be put in terms of z

OpenStudy (phi):

if I did the math correctly, we get 3x-3y+4z= (26/4)z -10/4 for z>=0 that means the smallest value is -10/4 and the largest is infinity

OpenStudy (qwertty123):

:)

OpenStudy (wcrmelissa2001):

:( thank you but the answer is actually -5/2≤3x-3y+4z≤67/7 which means it cant be infinity right? oh wait can it?

OpenStudy (wcrmelissa2001):

@phi i got the same answer as you as in x= ¼ z + ¼ y= -7/12 z +13/12 After which I put 1/4 z + 1/4 ≥0 z+1≥0 z≥-1 and -7/12 z + 13/12 ≥0 -7z+13≥0 z≥13/7 This doesn't make much sense to me actually since hten the answer would be z≥13/7. And this is only the lower limit, and i don't know how to find the upper limit

OpenStudy (youssefmohamed):

make one equation in one variable

OpenStudy (youssefmohamed):

If x,y,z≥0, 3y+2z=3+x, 3y+z=4-3x, find the largest and smallest value of 3x-3y+4z...\[x=3y+2z-3.(1)....x=(1/3)(4-3y-z)......from these two equations \]gety and z in terms of x and subistitute with them to in the form of equation in one variable

OpenStudy (wcrmelissa2001):

I know that's what I did. I made everything in terms of z. Afterwards, I used the z form of it to write: x≥0 (x in z form) and y≥0 (y in z form e.g. 2z) and solved it. But I wasn't able to get a range from this, I only managed to get one value. E.g. z≥2. This isn't a range and hence I'm unable to solve the question

OpenStudy (phi):

oh, you are correct! I just noticed that when we say x= ¼ z + ¼ y= -7/12 z +13/12 we also have to remember x>=0. The first equation is valid for all z>=0 but in the 2nd equation we need \[ y\ge0\\ -\frac{7}{12} z +\frac{13}{12} \ge 0 \\ \frac{13}{12} \ge \frac{7}{12} z \\ z \le \frac{13}{7} \] so we have an upper bound on z (else y would be negative) using that upper bound for z (not infinity!) we get \[ 3x-3y+4z= \frac{13}{2} z -\frac{5}{2} \\= \frac{13}{2}\cdot \frac{13}{7}-\frac{5}{2}\\ = \frac{67}{7} \] as the upper bound

OpenStudy (phi):

so the final result is \[ -\frac{5}{2} \le 3x-3y+4z\le \frac{67}{7} \]

OpenStudy (phi):

*** -7/12 z + 13/12 ≥0 -7z+13≥0 z≥13/7 **** I think that is where you ran into trouble. you have to "switch" the relation operator > to < if you multiply or divide by a negative number. However, it is always safe to add/subtract values from both sides. For example, you could start with -7z+13≥0 and add +7z to both sides to get 13 ≥ 7z 13/7 ≥ z which is the proper relation.

OpenStudy (phi):

and just to complete this, when you found *** 1/4 z + 1/4 ≥0 z+1≥0 z≥-1 **** we are told z ≥0, so we have to use 0 as at the lower bound on z, not -1

OpenStudy (wcrmelissa2001):

THANK YOU!! :D